# EmoDubber: Towards High Quality and Emotion Controllable Movie Dubbing

### v1

@@ 3.2 Lip-related Prosody Aligning
where $\mathcal{O}_{s}\in\mathbb{R}^{{P}\times{d_{m}}}$, and the ${d_{m}}$ and ${P}$ represent the hidden dimension and length of the phoneme sequence, respectively. $S_{id}$ is speaker style embedding, which is extracted by pre-trained speaker encoder from reference audio $R_{a}$, following [7, 50]. Next, we adopt prosody adaptor [6, 50] to generate phoneme-level prosody variations: pitch embedding ${P}_{pho}$ and energy embedding ${E}_{pho}$. Thus, the phoneme-level prosody sequences $\mathcal{O}_{p}=\mathcal{O}_{s}\oplus{P}_{pho}\oplus{E}_{pho}$ is calculated by combining style phoneme embedding and prosody variations. To obtain the lip-motion embedding from the input video frame sequence $V_{f}$, we adopt the same extracting pipeline as [6]:
$\mathcal{E}=\mathrm{LipEncoder}(M_{roi}\in\mathbb{R}^{{T_{v}}\times{D_{w}}\times{D_{h}}\times{D_{c}}}),$ (3)
where $M_{roi}$ indicates the mouth Region of Interest (ROI) frame sequence cropped by face landmarks from $V_{f}$, following [6]. $D_{w}$, $D_{h}$, and $D_{c}$ indicate the number of width, height, and channels of images in the mouth ROI frame. $T_{v}$ denotes the total length of mouth ROI frame. $\mathcal{E}\in\mathbb{R}^{{T_{v}}\times{d_{m}}}$ denotes the output lip motion embedding.
@@
where $C_{pho}\in\mathbb{R}^{{{T_{v}}}\times{d_{m}}}$ denotes the lip-prosody context sequences with the same length with video clips. Instead of non-constraint [6, 22] or simple diagonal-constraint [7, 16], our LPA module focuses on duration-level contrastive learning (DLCL) to achieve corrected alignment to guarantee monotonicity and subjectivity of weight matrix between prosody and lip-motion sequences:
@@
where $M_{lip,pho}^{gt}$ is a “0-1” matrix with $P$-th row and $T_{v}$-th column and satisfies the monotonicity and surjectivity. The value “1” represents correct correspondence between textual phoneme and lip motion through Montreal Forced Aligner (MFA) [33] model and coefficient between Frames per Second of the video (FPS) and sampling rate (SR). In this case, the DLCL encourages the positive pair to have a higher similarity to ensure the phoneme prosody unit focuses on the strongly related part in lip motion sequence.
@@ 3.3 Pronunciation Enhancing
specifically, $D_{p}=\mathrm{MAS}(\mathcal{E},\mathcal{O}_{p})\in\mathbb{R}^{{P}\times{1}}$ denotes the explicit duration for each phoneme unit, which records the integer multiple alignment between phonemes and video frames. The $\mathrm{LR}{(\cdot})$ is a length regulator, which aims to expand style phoneme sequences $\mathcal{O}_{s}\in\mathbb{R}^{{P}\times{d_{m}}}$ to video-level phoneme enhancement sequences $\mathcal{O}_{s}^{v}\in\mathbb{R}^{{T_{v}}\times{d_{m}}}$.
@@
where $\mathcal{V}_{f}\in\mathbb{R}^{{T_{v}}\times{d_{m}}}$ indicates the fused intermediate feature. The $\mathrm{Conformer}(\cdot)$ represents AVEC, which consists of an early fusion (EF) strategy [30] to reduce model complexity, 5 Conformer blocks (CB) without downsampling, and connectionist temporal classification (CTC) [12, 21] layer to maximize the sum of probabilities of correct target phoneme to ensure pronunciation.
@@ 3.4 Speaker Identity Adapting
where $\mu\in\mathbb{R}^{{M_{l}}\times{d_{a}}}$, and the ${M_{l}}$ and $d_{a}$ represent the length and hidden dimension of desired mel-spectrogram sequence. The SIA consists of up-sampling layer, utterance-level style learning (USL) module [7], and projection layere [7]. Firstly, we upsample the time axis of $\mathcal{V}_{f}$ to mel-spectrogram level by applying two layers of 2D convolutions [49]. Then, we use USL to inject the style information from style embedding $S_{id}$. Finally, the projection layey is used to project the output feature to target dimension.
@@ 4.1 Implementation Details
Video frames are sampled at 25 FPS and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following [32, 29]. The window length, frame size, and hop length in STFT are 640, 1024, and 160, respectively. The flow-prediction network is pre-trained on LibriSpeech [38] using 2 downsampling blocks, 2 midblocks, and 2 upsampling blocks. Each block had one Transformer layer with hidden dimensionality 256 and style affine layers [7], 2 heads, attention dimensionality 64, and snakebeta [25] activations. We use 8 heads for multi-head attention in LPA with 256 hidden sizes. The temperature coefficient $\tau$ of $\mathcal{L}_{cl}$ as 0.1. There are 5 conformer blocks in AVEC. We trained V2C, Chem, and GRID with batch sizes 64, 32, and 64, respectively. The Emotion Expert Classifier $\psi$ is trained on 13 emotional datasets with more than 50,000 emotional audio recordings in Emobox [31], which collects large-scale Speech Emotion Recognition (SER) benchmarks. During the inference process, $\gamma$ is set to 15, while $\alpha\in[0,5]$ and $\beta\in[0,2]$. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU. More details about the model and training are given in the supplementary material.
@@ 4.2 Datasets
GRID is a dubbing benchmark for multi-speaker dubbing [8]. The whole dataset has 33 speakers, each with 1000 short English samples. All participants are recorded in studio with unified background. The number of train and test data are 32,670 and 3280, respectively.

### v2

@@ 3.2 Lip-related Prosody Aligning
where $\mathcal{O}_{s}\in\mathbb{R}^{{P}\times{d_{m}}}$, and the ${d_{m}}$ and ${P}$ represent the hidden dimension and length of the phoneme sequence, respectively. $S_{id}$ is speaker style embedding, which is extracted by pre-trained speaker encoder from reference audio $R_{a}$, following [7, 50]. Next, we adopt prosody adaptor [6, 50] to generate phoneme-level prosody variations: pitch embedding ${P}_{pho}$ and energy embedding ${E}_{pho}$. Thus, the phoneme-level prosody sequences $\mathcal{O}_{p}=\mathcal{O}_{s}\oplus{P}_{pho}\oplus{E}_{pho}$ is calculated by combining style phoneme embedding and prosody variations. To obtain the lip-motion embedding from the input silent video clip $V_{l}$, we adopt the same extracting pipeline as [6]:
$\mathcal{E}=\mathrm{LipEncoder}(M_{roi}\in\mathbb{R}^{{F}\times{D_{w}}\times{D_{h}}\times{D_{c}}}),$ (3)
where $M_{roi}$ indicates the mouth Region of Interest (ROI) frame sequence cropped by face landmarks from $V_{l}$, following [6]. $D_{w}$, $D_{h}$, and $D_{c}$ indicate the number of width, height, and channels of images in the mouth ROI frame. $F$ denotes the total length of mouth ROI frame. $\mathcal{E}\in\mathbb{R}^{{F}\times{d_{m}}}$ denotes the output lip motion embedding.
@@
where $C_{pho}\in\mathbb{R}^{{{F}}\times{d_{m}}}$ denotes the lip-prosody context sequences with the same length with video clips. Instead of non-constraint [6, 22] or simple diagonal-constraint [7, 16], our LPA module focuses on duration-level contrastive learning (DLCL) to achieve corrected alignment to guarantee monotonicity and subjectivity of weight matrix between prosody and lip-motion sequences:
@@
where $M_{lip,pho}^{gt}$ is a “0-1” matrix with $P$-th row and $F$-th column and satisfies the monotonicity and surjectivity. The value “1” represents correct correspondence between textual phoneme and lip motion through Montreal Forced Aligner (MFA) [33] model and coefficient between Frames per Second of the video (FPS) and sampling rate (SR). In this case, the DLCL encourages the positive pair to have a higher similarity to ensure the phoneme prosody unit focuses on the strongly related part in lip motion sequence.
@@ 3.3 Pronunciation Enhancing
specifically, $D_{p}=\mathrm{MAS}(\mathcal{E},\mathcal{O}_{p})\in\mathbb{R}^{{P}\times{1}}$ denotes the explicit duration for each phoneme unit, which records the integer multiple alignment between phonemes and video frames. The $\mathrm{LR}{(\cdot})$ is a length regulator, which aims to expand style phoneme sequences $\mathcal{O}_{s}\in\mathbb{R}^{{P}\times{d_{m}}}$ to video-level phoneme enhancement sequences $\mathcal{O}_{s}^{v}\in\mathbb{R}^{{F}\times{d_{m}}}$.
@@
where $\mathcal{V}_{f}\in\mathbb{R}^{{F}\times{d_{m}}}$ indicates the fused intermediate feature. The $\mathrm{Conformer}(\cdot)$ represents AVEC, which consists of an early fusion (EF) strategy [30] to reduce model complexity, 5 Conformer blocks (CB) without downsampling, and connectionist temporal classification (CTC) [12, 21] layer to maximize the sum of probabilities of correct target phoneme to ensure pronunciation.
@@ 3.4 Speaker Identity Adapting
where $\mu\in\mathbb{R}^{{M_{l}}\times{d_{a}}}$, and the ${M_{l}}$ and $d_{a}$ represent the length and hidden dimension of desired mel-spectrogram sequence. The SIA consists of up-sampling layer, utterance-level style learning (USL) module [7], and projection layer [7]. Firstly, we upsample the time axis of $\mathcal{V}_{f}$ to mel-spectrogram level by applying two layers of 2D convolutions [49]. Then, we use USL to inject the style information from style embedding $S_{id}$. Finally, the projection layer is used to project the output feature to target dimension.
@@ 4.1 Implementation Details
Video frames are sampled at 25 FPS and all audios are resampled to 16kHz. The lip region is resized to 96 $\times$ 96 and pre-trained on ResNet-18, following [32, 29]. The window length, frame size, and hop length in STFT are 640, 1,024, and 160, respectively. The flow-prediction network is pre-trained on LibriSpeech [38] using 2 downsampling blocks, 2 midblocks, and 2 upsampling blocks. Each block had one Transformer layer with hidden dimensionality 256 and style affine layers [7], 2 heads, attention dimensionality 64, and snakebeta [25] activations. We use 8 heads for multi-head attention in LPA with 256 hidden sizes. The temperature coefficient $\tau$ of $\mathcal{L}_{cl}$ as 0.1. There are 5 conformer blocks in AVEC. We trained V2C, Chem, and GRID with batch sizes 64, 32, and 64, respectively. The Emotion Expert Classifier $\psi$ is trained on 13 emotional datasets with more than 50,000 emotional audio recordings in Emobox [31], which collects large-scale Speech Emotion Recognition (SER) benchmarks. During the inference process, $\gamma$ is set to 15, while $\alpha\in[0,5]$ and $\beta\in[0,2]$. Both training and inference are implemented with PyTorch on a GeForce RTX 4090 GPU. More details about the model and training are given in the supplementary material.
@@ 4.2 Datasets
GRID is a dubbing benchmark for multi-speaker dubbing [8]. The whole dataset has 33 speakers, each with 1,000 short English samples. All participants are recorded in studio with unified background. The number of train and test data are 32,670 and 3,280, respectively.